# MonoSOWA: Scalable monocular 3D Object detector Without human Annotations

### v1

@@ III-A Auto-labelling pipeline
Figure 4: Temporal aggregation examples. The point cloud of a significantly occluded vehicle (a) is recovered in subsequent frames (b) where the vehicle is no longer occluded. Distant car with poor depth prediction (c) is refined as the ego-vehicle drives closer (d). White points is pseudo-LiDAR of each frame, colourful points are the aggregated points. Human ground truth overlaid as green boxes for reference purposes.
Orientation and size estimation
Finding oriented 3D bounding boxes is a 7-degrees of freedom (DOF) problem. Instead of solving it directly, our method divides it into three disjunctive problems: orientation, location and dimensions.
@@
It is worth noting again that all locations are transformed into the reference coordinate space. To make it more robust, yaws for up to 5 locations before and after the reference frame are predicted, and then the median is taken out of those yaw predictions.
For stationary instances, we adapt the 3D box fitting algorithm of Zhang et al. [37]. The algorithm first flattens all points into BEV, and then it iterates over all possible angles $\theta\in{0,\frac{\pi}{2}}$. For each angle, two perpendicular axes to each other are computed, and all points are projected into both axes $C_{1}$ and $C_{2}$. Further, the criterion is computed, and $\theta$ with the minimal loss is chosen.
We propose a novel Saturated Closeness Criterion (algorithm 1). Instead of taking a simple sum of distances $d$, as Closeness Criterion in Zhang et al. [37], we feed the $d$ into a sigmoid multiplied by steepness parameter $\alpha$ (eq. 10), inspired by Template Fitting Loss in [31], to saturate the distance of outliers and reduce their influence on the fitting. As a result, the robustness and accuracy of estimation in noisy point clouds are significantly improved over the original algorithm.
$d$ $=\sigma(\alpha\cdot d)$ (10)
Another modification is that Closeness Criterion computes the distance to either minimum or maximum. However, those extremal points tend to be outliers. To mitigate the problem, each point is assigned either to the 10th or 90th percentile of the projected points instead of min/max. Intuitively, this creates an axis at the most probable location, mitigating outliers’ influence.
@@
1:$k$triangleright Steepness parameter
2:function SaturatedCloseness($C_{1},C_{2}$)
3:     $c_{1}^{\text{max}}\leftarrow P_{90}(C_{1})$, $c_{1}^{\text{min}}\leftarrow P_{10}(C_{1})$triangleright P is percentile
4:     $c_{2}^{\text{max}}\leftarrow P_{90}(C_{2})$, $c_{2}^{\text{min}}\leftarrow P_{10}(C_{2})$
5:     $D_{1}\leftarrow\arg\min_{v\in\{c_{1}^{\text{max}}-C_{1},C_{1}-c_{1}^{\text{min}}\}}\|v\|_{2}$
6:     $D_{2}\leftarrow\arg\min_{v\in\{c_{2}^{\text{max}}-C_{2},C_{2}-c_{2}^{\text{min}}\}}\|v\|_{2}$
7:     $D_{1}\leftarrow\sigma(\alpha\cdot D_{1})$triangleright $\sigma$ is the sigmoid function
8:     $D_{2}\leftarrow\sigma(\alpha\cdot D_{2})$
9:     $L\leftarrow 0$
10:     for  $i=1$ to $\text{length}(D_{1})$  do
11:         $d\leftarrow\min(D_{1}(i),D_{2}(i))$
12:         $L\leftarrow L+d$
@@
As mentioned, the algorithm chooses the yaw with a minimal loss. It creates two hypotheses, as it doesn’t differentiate between the front and back of the car. It is worth noting that the novel Saturated Closeness Criterion does not significantly increase the computational requirements.
To make the estimation more robust, we take both the $A_{j}$ and multiple $P_{i,j}$. Our method chooses frames so the object instance is the closest possible to the ego-vehicle, but at the same time, it is still fully in view to benefit from the fact that the pseudo-lidar is the most accurate for close predictions.
@@
The algorithm outputs a single spatial dimension for a single frame. If any of the values exceed the expected dimensions of an instance, the output is replaced by a prior estimate of the dimensions of a generic instance. Also, it is necessary to take into account that estimating the spatial dimensions of an instance from a single image can be an ill-posed problem; for example, if an instance is seen from the back, it is not possible to estimate length correctly. Thus, such cases are detected, and again, the output is replaced by a prior car dimensions estimate.
Position refinement
@@
The yaw $\theta_{i,j}$ and an approximate 3D position $L_{i,j}$ are already known from the previous steps. In order to get a more fine-grained estimate of position $L_{i,j}$, we apply small perturbations (up to 2 meters) along $x$ and $z$ axes and use Template Fitting Loss (TFL) [31] as the criterion to select the final position. In this stage, we also determine whether the car is facing towards or away from the ego-vehicle by testing two hypotheses for vehicle orientation – $\theta$ and $\theta+\pi$ and again selecting the orientation with the lowest TFL. As the orientation, location and dimensions are known at this step, 3D pseudo labels are created and further used to train the Monocular 3D Detector.
III-B Canonical Object Space (COS)
Figure 5: Inherent ambiguity caused by different camera focal length. The same scene (left) captured by two cameras with different focal length (middle) results in two images (right) where the object has different size. A monocular 3D object detector is however expected to predict the same 3D object location for both images, which is impossible without the detector knowing the focal length as well.
Camera focal length is a crucial parameter for precise metric depth estimation, but monocular 3D detectors, however, have no direct way to compensate for changing focal length because the same object at the same distance from the camera will have different size in the image, depending on the focal length (see fig. 5).
Assuming an object of size $S^{\text{real}}$ with a distance $D$ from the camera in world coordinates, the apparent size ($\approx$ number of pixels) in the image $S^{\text{img}}$ is given as
$\frac{S^{\text{real}}\cdot f_{1}}{D}=S^{\text{img}}_{1}\neq S^{\text{img}}_{2}=\frac{S^{\text{real}}\cdot f_{2}}{D}$ (11)
where $f_{1}$ and $f_{2}$ denote the focal length of camera capturing the same scene.
Taken to the extreme, for the exact same object image (same size $S^{\text{img}}$), the network needs to predict arbitrarily different distances in world coordinates due to the focal length change. This is clearly impossible, and this phenomenon confuses the network both in training as well as at inference time. When combining multiple datasets that were not taken with a single camera or even a single dataset that used different cameras to capture the dataset [38], it is a pre-requisite to be able to accommodate focal length differences.
Figure 6: Canonical Object Space. One scene viewed with two different cameras with different parameters. As the imaging size of the same object is different, 3D labels are scaled accordingly to adapt for focal length change. The predictions and loss computations are all calculated in the Canonical Object Space, and the final predictions are obtained after scaling back based on (known) camera parameters.
Inspired by Metric3D [3] where depth maps are normalized into a uniform space, our method presents Canonical Object Space (COS) where3D bounding boxes are trained to be invariant to changes in the focal length by choosing a single canonical focal length and then transforming the labels (objects) into COS. Instead of transforming whole images or depth maps, as in Metric3D, only the $x$,$y$,$z$ coordinates of all pseudo-ground truth labels are transformed. To transform labels, the scaling parameter $\omega_{i}$ is calculated as
$\omega_{i}=\frac{f^{C}}{f_{i}}$ (12)
$x^{C}=x\cdot\omega_{i}\qquad y^{C}=y\cdot\omega_{i}\qquad z^{C}=y\cdot\omega_{i}$ (13)
The model then operates in the COS and is directly supervised by the transformed labels (objects). During inference, all predictions are transformed from COS by dividing the depth prediction by scaling parameter $\omega_{j}$ of the inference frame $j$
$x=\frac{x^{C}}{\omega_{j}}\qquad y=\frac{y^{C}}{\omega_{j}}\qquad z=\frac{z^{C}}{\omega_{j}}$ (14)
This very simple transformation enables the network to train on multiple datasets where focal length differs and also allows the network to work in unseen camera setup (focal length) because $\omega_{j}$ is calculated at inference time (see table III – 0% fine-tuning row).
IV Experiments
Datasets
We use two public datasets for the evaluation – KITTI [8] and KITTI-360 [39]. KITTI dataset setup has four front-facing cameras, IMU and LiDAR. KITTI-360 setup uses two front-facing cameras and two fish-eye cameras facing sides, IMU and LiDAR. The cameras and their parameters differ in both datasets. For the KITTI dataset, the same training (3712 samples) and validation (3769 samples) splits as [2, 40, 29, 31] are used. For the KITTI-360 dataset, the training (6 sequences, 44178 frames), validation (2 sequences, 1115 frames) and testing (1 sequence, 2459 frames) splits from VSRD [4] are used. It is worth noting that, unlike VSRD, we are able to use all frames of the training set as our method is much faster.
Implementation
We used off-the-shelf MViT2-Huge [41] in the Detectron2 framework [42] trained on MS-COCO [43] as the 2D object detector and Metric3DV2-giant [9] for metric depth estimation. We use our pseudo-labels to train MonoDETR [19] monocular 3D object detector, using AdamW [44] optimiser with learning rate and weight decay equal to 0.0002 and 0.0001, respectively, while keeping other hyper-parameters as in  [19]. The canonical focal length is equal to 500. Aggregation is done over 100 frames, the $p$-score threshold for stationary/moving classification is 0.0001, and the minimum threshold for $d_{real}$ is 5 meters. The steepness parameter $\alpha$ in the Saturated Closeness Criterion is 10.
IV-A Results
@@
TABLE I: Evaluation results of monocular 3D object (vehicle) detection on the KITTI-360 test set. * denotes that given data modality was not used by the method itself, but it was required in the annotation process to create human 3D box annotations.
@@
TABLE II: Monocular 3D object detection accuracy on the KITTI validation set, depending on which datasets/labels are used in training.
KITTI-360.
The first evaluation and comparison with prior work is on KITTI-360 [39] as shown in table I. Our method shows outstanding results on both the Bird’s eye view (BEV) and 3D at 0.5 IoU as it outperforms current state-of-the-art weakly-supervised method VSRD [4], despite the fact our method does not use any human labels.
In the 0.3 IoU metric both for BEV and 3D, our method outperforms both WeakM3D [1] and Autolabels [2] by a significant margin and it achieves a competitive AP as VSRD [4]. We believe that this is likely caused by the detection failures, class ambiguities or that KITTI 2D bounding boxes are amodal as using consistent ground-truth 2D labels significantly helps VSRD.
Cross-dataset evaluation
Thanks to the proposed Canonical Object Space, it is possible to train on multiple merged datasets. We, therefore, combine both KITTI and KITTI-360 datasets training subsets, generate our pseudo-labels, train a single model and evaluate it on the KITTI validation subset (see table II). We show that generating pseudo-labels for two different datasets and then performing training on the joint significantly improves average precision both for 0.3 and 0.5 IoU on BEV and 3D. The gain comes from the additional amount of data present in the KITTI-360 dataset.
We also demonstrate the ability of our method to work out-of-the-box in previously unseen camera setup (see table II - row K360), where the model was trained using pseudo-labels on KITTI-360 but evaluated on in this case unseen KITTI dataset. We note that considering IoU 0.3, the accuracy of a model trained on KITTI-360 is actually even slightly higher than the model trained on the original domain (KITTI), which we contribute simply to the fact that KITTI-360 contains more training data.
Semi-supervised
GT APBEV/AP3D@0.7
Method Ratio Easy Moderate Hard
VSRD [4] 0% 0.002/0.001 0.004/0.001 0.005/0.002
Ours 24.46/8.24 20.03/6.84 16.14/6.36
VSRD [4] 25% 31.72/21.76 22.32/15.43 18.86/12.55
Ours 39.99/32.64 30.92/25.06 26.47/20.35
VSRD [4] 50% 43.44/31.05 31.54/21.48 27.17/17.93
Ours 44.32/33.70 32.97/25.65 27.52/23.08
VSRD [4] 75% 42.58/32.95 31.08/24.68 27.19/21.38
Ours 43.32/34.11 32.48/25.68 27.39/23.05
MonoDETR [19] 100% 37.99/29.36 26.76/20.64 23.02/17.30
TABLE III: 3D object detection accuracy on KITTI validation set depending on the fraction of human labels used to fine-tune the model.
In this experiment, MonoDETR is trained jointly on KITTI and KITTI-360, while all labels are generated by our auto-labelling pipeline, as a pre-training step. Next, only a given fraction of human labels is used to fine-tune the model.
Firstly, as seen in table III, thanks to COS our method works out-of-the-box without any fine-tuning (0% human labels), unlike previous methods, showing there is no need to fine-tune using human labels while moving to different domains. Secondly, pre-training on a larger dataset auto-labelled by our method and then fine-tuning on the target domain significantly improves the average precision over simply training on the target domain from scratch. With only 25% of the ground truth labels, our method outperforms traditional fully-supervised detector trained from scratch with 100% of human labels, and the gap grows as more human labels are used. Lastly, our method significantly outperforms current state-of-the-art VSRD [4] by a large margin, showing that pre-training on larger datasets is a promising direction for further research.
Figure 7: Example 3D car detections of our model trained without using human annotations on both KITTI [8] and KITTI-360 [39].

### v2

@@ III-B Local Object Motion Model
Figure 4: The point cloud of a significantly occluded vehicle (a) is recovered in subsequent frames (b). Distant car with poor depth prediction (c) is refined as the ego-vehicle drives closer (d). Pseudo-LiDAR of each frame in white, aggregated object points in purple, human label in green for reference.
III-C Step-wise autolabelling
Finding oriented 3D bounding boxes is a 7-degrees of freedom (DOF) problem. Instead of solving it directly, our method divides it into three disjunctive problems: orientation, location and size estimation.
Orientation and size estimation.
@@ III-C Step-wise autolabelling
To make it more robust, yaws for up to 5 locations before and after the reference frame are predicted, and then the median is taken out of those yaw predictions.
For stationary instances, we improve the 3D box fitting algorithm of Zhang et al. [41]. The algorithm first flattens all points into Bird’s eye view (BEV), and then it iterates over all possible angles $\theta\in{0,\frac{\pi}{2}}$. For each angle, two perpendicular axes to each other are computed, and all points are projected into both axes $x\in\mathbb{R}$ and $y\in\mathbb{R}$.
@@
1:$\alpha$triangleright Steepness parameter
2:function SaturatedCloseness($x,y$)
3:    $x^{\text{max}}\leftarrow P_{90}(x)$, $x^{\text{min}}\leftarrow P_{10}(x)$triangleright P is percentile
4:    $y^{\text{max}}\leftarrow P_{90}(y)$, $y^{\text{min}}\leftarrow P_{10}(y)$
5:    $x\leftarrow\arg\min_{v\in\{x^{\text{max}}-x,x-x^{\text{min}}\}}\|v\|_{2}$
6:    $y\leftarrow\arg\min_{v\in\{y^{\text{max}}-y,y-y^{\text{min}}\}}\|v\|_{2}$
7:    $x\leftarrow\sigma(\alpha\cdot x)$triangleright $\sigma$ is the sigmoid function
8:    $y\leftarrow\sigma(\alpha\cdot y)$
9:    $\mathcal{L}\leftarrow 0$
10:    for  $i=1$ to $\text{length}(x)$  do
11:        $e\leftarrow\min(x(i),y(i))$
12:        $\mathcal{L}\leftarrow\mathcal{L}+e$
@@
For each $\theta$, we calculate a newly proposed Saturated Closeness Criterion (algorithm 1). Instead of taking a simple sum of distances $x$ and $y$, as Closeness Criterion in Zhang et al. [41], we feed the $x$ and $y$ into a sigmoid multiplied by steepness parameter $\alpha$ (Line 6 and 7 of algorithm 1), inspired by Template Fitting Loss in [30], to saturate the distance of outliers and reduce their influence on the fitting. As a result, the robustness and accuracy of estimation in noisy point clouds are significantly improved over the original algorithm.
Another modification is that Closeness Criterion computes the distance to either minimum or maximum. However, those extremal points tend to be outliers. To mitigate the problem, each point is assigned either to the 10th or 90th percentile of the projected points instead of min/max. The assignment is based on the minimal distance to one of the extremal points. Intuitively, this creates an axis at the most probable location, mitigating outliers’ influence. The algorithm chooses the yaw with a minimal criterion value and creates two hypotheses, as it doesn’t differentiate between the front and back of the car.
@@
The algorithm also outputs a spatial dimensions estimate for a single frame. If any of the values exceed the typical dimensions of an instance, the output is replaced by a prior estimate of the dimensions of a generic instance. Also, it is necessary to take into account that estimating the spatial dimensions of an instance from a single image can be an ill-posed problem; for example, if an instance is seen from the back, it is not possible to estimate length correctly. Thus, such cases are detected by computing the difference between car orientation and the viewing angle of the car and if the difference is ${0,\frac{\pi}{2},\pi,\frac{3\pi}{2}}$, then the output is replaced by a prior car dimensions estimate.
Position refinement.
@@
The yaw $\theta$ and a prior 3D position are already known from the previous steps. In order to get a more fine-grained estimate of position, we apply small perturbations (up to 2 meters) along $x$ and $z$ axes and use Template Fitting Loss (TFL) [30] as the criterion to select the final position. To compute TFL, we use a generic template mesh of a vehicle sampled into a point cloud. Given a generic vehicle point cloud, we move the point cloud according to perturbations and compute TFL at those points. The lowest TFL value corresponds to the best fit between those point clouds. Unlike Chamfer Distance, it works well with outliers. In this stage, we also determine whether the car is facing towards or away from the ego-vehicle by testing two hypotheses for vehicle orientation – $\theta$ and $\theta+\pi$ and again selecting the orientation with the lowest TFL.
Figure 5: Generated 3D labels are transformed into Canonical Object Space to accommodate for different focal lengths; the resulting model is therefore trained to output canonical coordinates. At inference time, model predictions are transformed back to world coordinates, using (known) camera parameters.
@@
TABLE I: Evaluation results of monocular 3D object (vehicle) detection on the KITTI-360 [42] test set. * denotes that given data modality was not used by the method itself, but it was required in the annotation process to create human 3D box annotations.
III-D Canonical Object Space (COS)
Camera focal length is a crucial parameter for precise 3D position estimation, but monocular 3D detectors have no direct way to compensate for changing focal length because the same object at the same distance from the camera will have different size in the image, depending on the focal length.
Taken to the extreme, for the exact same object image measured in pixels, the network needs to predict arbitrarily different distances in world coordinates, depending on camera focal length. This phenomenon confuses the network during training, and it is therefore necessary to solve this problem in order to be able to train a network on data captured by different cameras.
Inspired by methods such as Metric3D [5] and Omni3D [6], our auto-labelling method uses Canonical Object Space (COS) where 3D bounding boxes are trained to be invariant to changes in the focal length by choosing a single canonical focal length and then transforming the generated labels (objects) into COS. Instead of transforming whole images or depth maps as in Metric3D, only the $x,y,z$ coordinates of pseudo-ground truth labels are transformed directly when creating the pseudo-labels. To transform labels, the scaling parameter $\omega_{i}$ is calculated as
$\omega_{i}=\frac{f^{C}}{f_{i}},$ (8)
$x^{C}=x\cdot\omega_{i}\qquad y^{C}=y\cdot\omega_{i}\qquad z^{C}=y\cdot\omega_{i}.$ (9)
The model then operates in the COS and is directly supervised by the transformed labels (objects). During inference, all predictions are transformed from COS using the scaling parameter $\omega_{j}$ of the inference frame $j$ (see fig. 5).
This very simple transformation enables the network to train on multiple datasets where focal length differs and also allows the network to work in unseen camera setup (focal length) because $\omega_{j}$ is calculated during inference.
IV Experiments
APBEV/AP3D@0.5
Method Published Easy Hard
Autolabels [2] ICCV2021 51.85/4.65 46.10/2.92
VSRD [4] CVPR2024 47.12/35.25 43.91/32.64
Ours - 61.17/47.07 51.92/45.51
TABLE II: Auto-labelling AP on KITTI-360 [42] training set.
Training supervision 3D Object detection accuracy (Level 2)
Human Labels APBEV@0.5 AP3D@0.5
Method LiDAR Masks 3D boxes All 0m-30m 30m-50m 50m-$\infty$ All 0m-30m 30-50m 50m-$\infty$
DEVIANT [43] yes* yes 10.29 26.75 4.95 0.16
MonoDETR [20] yes* yes 23.63 39.10 20.95 4.70 21.41 37.89 18.61 3.69
MonoSOWA (Ours) 18.98 33.51 12.02 4.55 13.46 24.65 5.87 4.55
TABLE III: Monocular 3D object (vehicle) detection AP on Waymo [44] validation set.
Datasets.
We use three public datasets for evaluation – KITTI [10], KITTI-360 [42] and Waymo [44]. In the KITTI dataset, we employ the right RGB camera; in the KITTI-360 dataset, we use the right perspective camera; and in Waymo, we utilize the front-facing camera. Note that the camera parameters – their focal length, resolution, and FOV – are different in each dataset. In KIITI, we use the same training (3,712 samples) and validation (3,769 samples) splits as [2, 45, 29, 30] are used. In KITTI-360 dataset, training (6 sequences, 44,178 frames) and testing (1 sequence, 2,459 frames) splits from VSRD [4] are used. Waymo experiments follow the official training (798 sequences, 158,080 frames) and validation (202 sequences, 39,988 frames) splits.
Implementation.
We used off-the-shelf MViT2-Huge [40] in the Detectron2 framework [46] trained on MS-COCO [47] as the 2D object detector and Metric3DV2-giant [11] for metric depth estimation. We use our pseudo-labels to train MonoDETR [20] monocular 3D object detector, using AdamW [48] optimiser with learning rate and weight decay equal to 0.0002 and 0.0001, respectively, while keeping other hyper-parameters as in  [20]. The canonical focal length $f_{c}$ is equal to 750. Aggregation is done over 100 frames, the $T_{z}$ threshold for stationary/moving classification is 0.2, and the minimum threshold $T_{m}$ is 5 meters. The steepness parameter $\alpha$ in the Saturated Closeness Criterion is 10 (see Supplementary material for ablations). Note that we used the same hyper-parameter settings for all three datasets.
IV-A Results
Training Fine-tuning APBEV/AP3D@0.7
pseudo labels human labels Easy Moderate Hard
VSRD [4] 0% 0.002/0.001 0.004/0.001 0.005/0.002
Ours 24.46/8.24 20.03/6.84 16.14/6.36
VSRD 25% 31.72/21.76 22.32/15.43 18.86/12.55
Ours 39.99/32.64 30.92/25.06 26.47/20.35
VSRD 50% 43.44/31.05 31.54/21.48 27.17/17.93
Ours 44.32/33.70 32.97/25.65 27.52/23.08
VSRD 75% 42.58/32.95 31.08/24.68 27.19/21.38
Ours 43.32/34.11 32.48/25.68 27.39/23.05
- 100% 37.99/29.36 26.76/20.64 23.02/17.30
TABLE IV: Monocular 3D object detection evaluated on KITTI [10] validation set. Pre-training on KITTI-360 [42] pseudo-labels and fine-tuning on a fraction of the KITTI training set.
KITTI-360.
In the first experiment, we generate pseudo-labels on the KITTI-360 [42] dataset and train MonoDETR [20] with these pseudo-labels. We then compare our trained model with previously reported results of other trained monocular 3D object detectors. As shown in table I, our method outperforms current state-of-the-art weakly-supervised method VSRD [4] both in Bird’s eye view (BEV) and 3D average precision at 0.5 IoU, despite the fact our method does not use any human labels. In the 0.3 IoU metric, our method outperforms both WeakM3D [1] and Autolabels [2] by a significant margin both for BEV and 3D and it achieves competitive results with VSRD [4]. The remaining gap in the IoU 0.3 metric is caused by KITTI-360 human labels which are amodal (they include occluded or invisible object parts), which significantly helps VSRD that uses them in its training. Our method instead relies on a 2D detector, which only outputs boxes based on what is actually visible in the image, but as shown in table X, when identical inputs are used, our method significantly outperforms VSRD in all metrics. Note that our auto-labelling pipeline takes, on average, 1.3 seconds per frame (including all the models it uses as inputs - see table IX) – a 700x speed-up.
We also directly compare all generated labels with human labels on the KITTI-360 training set (see table II). We show significantly improved accuracy of our pseudo-labels compared to previous auto-labelling methods [4, 2], both in 3D and BEV.
Waymo.
In the second experiment, we evaluate our method on Waymo dataset [44], where we again train MonoDETR [20] detector using only our pseudo-labels generated for Waymo training set and evaluate on Waymo validation set (see table III). Despite using no human annotations, our method achieves competitive performance to methods using human annotations. Note that we are unable to directly compare to VSRD [4] on this dataset, because the training set has 150k frames, and it is therefore practically infeasible to auto-label the data using VSRD (15 minutes per frame) in any reasonable time frame.
Pseudo-labels for pre-training.
In the following experiments, the labels generated by our auto-labelling pipeline are used as a pre-training step and then only a small fraction of human labels is used to further fine-tune the model.
Training Fine-tuning APBEV AP3D@0.5
pseudo labels human labels
Ours 1% 21.59 18.43
- 1% 15.55 13.81
Ours 5% 22.56 20.68
- 5% 21.08 18.12
Ours 10% 22.80 20.65
- 10% 21.69 18.53
Ours 15% 23.20 20.88
- 15% 22.63 20.60
- 100% 23.63 21.41
TABLE V: Monocular 3D object detection (vehicle - Level 2) on Waymo validation set. Training on Waymo training set using pseudo-labels and fine-tuning on a fraction of human labels.
First, in table IV we show that pre-training on a larger dataset (KITTI-360) auto-labelled by our method and then fine-tuning on the target domain (KITTI) significantly improves the average precision when compared to simply training on the target domain from scratch. With only 25% of human labels, our method outperforms traditional fully-supervised detector trained from scratch with 100% of human labels, and the gap grows as more human labels are used. Also note that our method significantly outperforms current state-of-the-art VSRD [4] by a large margin.
Second, we show on a large-scale dataset (Waymo) that pre-training with our method can save human annotations costs. In this scenario, we assume the annotation budget is limited and therefore it’s not possible to label all images in the training set. We show, that when only a fraction of images are labelled by a human – e.g. only 5%, using our pseudo-labels as a pre-training step leads to consistent improvements in detection accuracy at no annotation cost, compared to only using available human labels (see table V). Also, when using only 15% of human labels and our method as a pre-training step, the accuracy is almost identical to using all human labels; this can also be viewed that using our method leads to saving 85% of annotation costs for a very small drop in accuracy.
@@ IV-A Results
TABLE VI: MonoDETR [20] AP on the KITTI [10] validation set, depending on which datasets/pseudo-labels are used in training.
Cross-dataset evaluation.
Thanks to using Canonical Object Space, it is possible to combine datasets with different camera setups. First, we demonstrate that unlike previous work [4], our method works out-of-the-box in previously unseen camera setup (see table VI - row K360 and table IV - row 0%), where the model was trained using pseudo-labels on KITTI-360 but evaluated on in this case unseen KITTI dataset. We also combine both KITTI and KITTI-360 datasets training subsets, generate our pseudo-labels, train a single model and evaluate it on the KITTI validation subset (table VI – row KITTI+K360). Combining pseudo-labels from KITTI+K360 achieves comparable accuracy to using human labels of KITTI, and even surpasses the accuracy of training with human labels in the Hard category in the 0.3 metric.
Figure 6: Example of 3D car detections by MonoDETR [20] trained without human annotations on KITTI [10] (top), KITTI-360 [42] (middle) and Waymo [44] (bottom)
Qualitative analysis.
We provide qualitative analysis in fig. 6 and in the Supplementary material in fig. 7 and fig. 8.